# A Non-Traditional Approach to Assisting Data Address Translation

### v1

@@ III-E Address bits for PCAT access
The bits used for a PCAT access could be from the PC of a (problem) load instruction, or the bits of the cache block number containing a (problem) load instruction. The former allows more flexibility as an arbitrary number of loads from the same cache block can reside in a PCAT. However, this requires a per-instruction APT bit in the L1i cache and also multiple accesses to the PCAT1 if these loads reside in different sets of the PCAT1. In the latter case, there needs to be only one APT bit, and one PCAT1 access per cache block. Multiple VPN-PTE entries for multiple problem loads in a cache block can reside in the different ways of a set and thus be obtained in parallel, with the set associativity capping the number of problem loads per cache block for which the PCAT1 can maintain information. We briefly evaluate the two options in section IV-E.
@@ III-F PCAX Coupling with the CDTLB/STLB
A question that arises is whether the structures for PCAX (PCAT1/PCAT2) and the conventional address translation (CDTLB/STLB) should be coupled or not. That is, when a PTE is brought into the PCAT1, as a result of a CDTLB/STLB miss, should the PTE also be placed in the CDTLB/STLB (the structures are coupled), or solely in the PCAT1 (the structures are not coupled). Not coupling the structures frees up capacity in the CDTLB/STLB which could be used for other translations whereas not placing a PTE in the CDTLB when it is brought into the PCAT1 results in misses for other proximate load instructions that need the same PTE. Such instructions can’t get the translation from the PCAT1 (because of different PCs) and it won’t be there in the CDTLB.
Our experimental observations suggest that PCAT1/CDTLB coupling is beneficial for most applications as many other load instructions also access a PTE (from a CDTLB) brought in by a CDTLB-missing load. Accordingly, when a PTE is placed in the PCAT1, it is also placed in the CDTLB. For the STLB, given that there is little reason to maintain inclusion between the STLB and the CDTLB/PCATs, we do not couple the PCATs and the STLB. This reduces unnecessary duplication of entries in the PCAT1/CDTLB and STLB, and unnecessary evictions from the STLB, leading to a lower SLTB miss rate, as we shall quantify in section IV-H.
III-G PCAT/CDTLB/STLB and Page Table Interactions
Like a conventional CDTLB, a PCAT operates like a cache of PTEs, albeit one that is accessed with a load PC rather than a data address. When PTEs are cached, appropriate actions need to be taken to ensure correctness. When an update is made in a page table entry, copies cached in CDTLBs must be invalidated using a TLB Shootdown [47, 36]. These changes include the PPN mapping, permissions, and potential updates to the Access bit (if the page replacement algorithm requires it).
Since the PCAT is accessed with a load PC, and not with a data virtual address, operations that rely on a data virtual address (if contemplated) are more involved. TLB shootdown operation typically uses the data virtual addresses to perform invalidations. We invalidate all the entries in the PCATs on any change to an entry in the page table (a PCAT Shootdown). Since a PCAT entry is created on a (first) CDTLB miss, where the page walker also updates the Access bit in the page table and gives the correct value to the CDTLB, the PCAT will also have the correct value of the Access bit. Any subsequent clearing of the Access bit in the page table (by the OS) would result in a PCAT Shootdown if needed. Thus having a PCAT does not impact the operation of the page replacement algorithm.
The Dirty bit is more involved since the first write to a page can happen after an entry has been created in the PCAT. Moreover, the same PTE could be present with different load PCs in the PCATs. If a PTE is brought into the CDTLB from the PCAT (with an unset Dirty bit), rather than from the STLB or the page table (with a set Dirty bit), it simply results in an additional ”store miss” to set the Dirty bit in the CDTLB which then frequently hits in the STLB (with the Dirty bit set), only proceeding to the page table on an STLB miss.
@@ III-H Periodically Invalidating the PCATs
The PCATs maintain VPN-PTEs associated with load instructions. The (valid) VPN-PTEs are always correct since any update to a PTE will result in the invalidation of the entire PCATs, as above. However, some performance (not correctness) pathologies can result because unproductive entries can be held on to in the PCATs longer than needed and reduce their effective size. These situations can arise when a load instruction with a PCAT entry becomes less active (or cold). One example we have observed is that when multiple entries for a code cache block are brought into the PCAT1 from the PCAT2, they include entries associated with cold load instructions which are not used in the PCAT1, but are considered to be accessed by the PCAT2 replacement algorithm, and thus not evicted from the PCAT2. Another example is when the number of entries per cache block is limited (e.g., by PCAT associativity), entries created by dynamically earlier instructions that become cold take up slots that later, more active instructions, could profitably use. To handle such situations created by dynamically cold code we can periodically clear the PCATs and re-populate them again. Such periodic invalidation and re-population have also been found to be useful in memory dependence predictors [14, 31], where store-load relationships that have been learnt are cleared to remove relationships that have become cold.
@@ III-I Selectiveness of Loads Using PCAT
PCATs are intended to assist CDTLBs by providing translations that might miss in the CDTLB. PCAT sizes can be kept small because the number of static loads that miss in a CDTLB are small. However, PCAX could also be used for a small number of other ”high value” loads from the set of loads that would normally hit in the CDTLB (whose number is, of course, much larger). Like the data of figure 2 for loads that miss in a CDTLB, we have also found that many times a dynamic instance of a load that hits in a CDTLB uses the same PTE as the last dynamic instance, as we shall see in section IV-D. Judiciously using PCAX for loads that might normally hit in a CDTLB, while maintaining reasonably small-sized PCATs, to improve translation latency and energy is left for future work.
@@ IV Evaluation
Processor Decoupled Front-end
Fetch Target queue 192 instr
@@
Decode queue 60 entry
Dispatch queue 60 entry
Branch target buffer 8K entries
Target cache 4K entries
Return Address stack 64 entries
Branch penalty 2 cycles (decode stage)
@@
L1 DTLB 64-entry, 4-way, 1-cycle, 4-entry MSHR
@@
Page Structure Caches 3-level Split PSC, 2-cycle; PML4: 2-entry, fully;
PDP: 4-entry, fully; PD: 32-entry, 4-way
@@ IV-A Benchmarks and Experimental Infrastructure
For our evaluation, we use ChampSim [18], a detailed simulator that simulates an out-of-order core with 6-wide fetch and 4-wide issue. Table I summarizes the parameters of the simulated microarchitecture. This version of ChampSim also models a realistic page table walker as found in x86 architectures. Key aspects modeled include: (i) variable latency cost of page table walks and (ii) caching of page walk references in data memory hierarchy as in [50]. Specifically, this models a 4-level page table, a hardware page table walker, and a 3-level split page structured cache. The page walker supports up to 4 concurrent STLB misses. All caches and TLBs use a normal LRU replacement policy. The default page size is 4KB and the default PCAX scheme employs a 64-entry, 2-way PCAT1, a 2K entry, 8-way PCAT2, and a 2-entry RPA Buffer. PCAT1 uses normal LRU while PCAT2 uses the slightly modified LRU replacement which was explained in Section III-C.
For benchmark programs, we use a set of 84 server traces from Qualcomm, which were released after modifications recently [17, 16]. These traces have been widely used in recent papers using the Champsim simulation infrastructure, and studying a variety of issues in instruction supply, memory hierarchies, and microarchitecture [50, 49, 6, 39, 28, 23] and more which can be found in [20]. Though the data footprint of these traces—in the range of 20MB-100MB—isn’t as large as others that have been used in studies of data address translation [26, 25, 21], the data referencing patterns result in a range of data TLB miss rates, from 1% to 31% with a 64-entry, 4-way set associative, data TLB, making them an apt choice for evaluating PCAX. All traces are run to completion with a warm-up of 10M instructions. Many benchmarks traces have 100M instructions, while some of the traces are smaller (30M instructions) [3].
@@ IV-B Storage Costs
An entry in the PCAT1 takes 11.6B (20-bit tag, 1 valid bit, 1 useful bit, 1 STLBM bit, 38-bit virtual page, and 4B for the PTE). A 64-entry PCAT1 takes about 744B. PCAT2 with 512, 1K, and 2K entries takes about 6KB, 12KB, and 24KB, respectively. There are 512 APT bits in the 32KB, 64B block, L1i cache, taking up 64B. There are 8192 APT bits in the 512KB, 64B block, L2 cache, taking up 1024B. The RPA buffer takes up 18B.
We use a configuration with a 64 entry PCAT1 and 2K PCAT2 which amounts to a total storage of 25.8KB. However, even 512 and 1K entry PCAT2 provide similar performance for many of the applications, as we show in Section IV-F.
@@ IV-C Results Overview
We start our evaluation by presenting some relevant characteristics of the benchmark programs in section IV-D. Then section IV-F evaluates the reduction in miss rates for primary DTLBs for different PCAT configurations. Section IV-G studies the effect on larger page sizes. Section IV-H evaluates the reduction in miss rates for secondary DTLBs for different PCAT configurations. Section IV-I presents the service sites for data address translation for a select benchmark set and then sections IV-J and IV-K evaluate the impact on CPI and energy, respectively. Lastly, Section IV-L studies the effect of invalidating the PCATs periodically.
The ordering of the benchmarks in the different figures is different for better visual presentation.
Figure 5: Total Data Pages Touched
Figure 6: CDTLB Load Miss Rate
@@
IV-D Benchmark Characteristics
In this section, we consider some key benchmark characteristics that are relevant to our evaluation. Figure 5 presents the number of 4KB data pages touched by the different applications (note log scale on Y-axis). The data footprint is not very large. However, the benchmarks also have large code footprints [6, 5, 12, 39, 38]; for an L1i of size 32KB, there are 45 benchmarks with an L1i MPKI between 0-30, 11 benchmarks with an L1i MPKI between 30-45, 21 benchmarks with an L1i MPKI between 45-60 and 7 benchmarks with an L1i MPKI $>$ 60. The large code footprint results in very frequent code movement from L2 to L1i, and coupled with the data referencing pattern collectively results in a significant number of CDTLB misses, shown in Figure 6 for CDTLB sizes of 64, 128, 256, and 1024 entries, with 4-way associativity in each case. For a 64-entry CDTLB, most of the applications have a miss rate greater than 5%. Even for a (single-level) 1024-entry CDTLB, several applications have a miss rate greater than 2%.
@@ IV-D Benchmark Characteristics
Next, we consider problem loads that account for 90% of the CDTLB misses. Figure 7 shows the total number of problem loads for 64 and 128 entry CDTLBs. The total number is of the order of a few hundred and less than 2000 in almost all cases  [^5] In some cases the number is larger for a 128-entry CDTLB than for a 64-entry one. This is because those applications have a high miss rate with a 64-entry CDTLB (e.g., 20%), which comes down quite significantly (e.g., to 10%) with a 128-entry CDTLB. Accounting for 90% of the misses bring in more problem PCs, which aren’t included in the 90% when the miss rate is higher.. These numbers in this figure give an idea of how large of a PCAT2 might work well.
@@
Since the PCAT1 needs to accommodate only problem loads that are resident in the L1 instruction cache, we consider this number in Figure 8. The figure presents the maximum number of problem loads resident in the 32KB L1 instruction cache at any time for 64-entry and 128-entry CDTLBs. This maximum number is in the small hundreds. The average number (not shown) is smaller. As we shall see, PCAT1 sizes that can accommodate a few tens of problem loads (e.g., 64) are adequate in most cases.
Figure 9: Loads Accessing same PTE
Figure 10: PTE Reuse Characteristics
Figure 11: PTE Rereference Distance
In Section II, we analyzed the load references that miss in the CDTLB on the same PTE as the previous dynamic missing instance. Figure 9 presents the percentage of all loads (which includes both hits and misses in the CDLTB) that access the same PTE as the previous dynamic instance of the load. We observe that this percentage is as high as 70-90% for many of the benchmarks. This result should not be surprising as many dynamic instances of a static load access proximate elements of a data structure, which are likely to be in proximate memory addresses, and thus mostly on the same page.
Figure 10 presents the reuse characteristics of PTEs, that is, after an initial load reference that brings a PTE into the CDTLB, what percentage of reuses of the PTE (either from another dynamic instance of the same static load, or from dynamic instances of other static loads) occur with $<$ 10, $<$ 50, $>$ 100, or $>$ 1000 dynamic load instructions. From the figure we see that a very large percentage of references (and reuses) to a PTE occur within fewer than 50 dynamic load references and a very small percentage are for $>$ 100 references. Further, we have observed, data not shown, that dynamic instances of an average of 2-8 other static loads reference a PTE brought into a CDTLB by a dynamic instance of a static load. These results should also not be surprising because we expect references to different fields of an element of a data structure (from different static loads), which are likely to be on the same page, to occur close to each other.
Figure 11 presents the re-reference distance for a PTE. That is, between the time a PTE is brought into the CDTLB by a dynamic instance of a static load, how many other PTEs are referenced, on average, until the time the same PTE is re-referenced and brought into the CDTLB again by another dynamic instance of the same static load. Data is presented for re-reference distances of greater than 100, 200, 500, and 1000 PTEs. As we see, many benchmarks have more than 50-60% of such missing PTE references with a re-reference distance greater than 100 and 200. Even the percentage of such misses with re-reference distances greater than 500 and 1000 is non-trivial (20-30%) for some of the benchmarks. Such large re-reference distances make it difficult to maintain a given PTE in a relatively small-sized CDTLB, leading to high miss rates.
To put it together, a dynamic instance of a static load references a given PTE, which is then (mostly) quickly followed by dynamic instances of other static loads referencing the same PTE. The load then gets replaced from the L1i and the PTE from the CDTLB (at different points of time). Later the same load is brought into the L1i again, and then it references the same PTE again. This is what provides the basis for PCAX: associate PTEs with a small subset of loads (which miss in the CDTLB) and bring a PTE associated with a load into the PCAT1 when the load is brought back into the L1i so that it can be used when the load references the PTE without having to incur a CDTLB miss. In other words, view the issue of CDTLB misses not from the viewpoint of the data reference stream, but from the viewpoint of the instruction reference stream which creates the data references that result in the CDTLB misses. If a PTE is associated with a load, it can be accessed with the load PC, not with the address of the data being accessed.
IV-E PCAT Indexing

### v2

@@ III-D Address bits for PCAT access
The bits used for a PCAT access could be from the PC of a (problem) load instruction, or the bits of the cache block number containing a (problem) load instruction. The former allows more flexibility as an arbitrary number of loads from the same cache block can reside in a PCAT. However, this requires a per-instruction access PCAT bits in the L1i cache and also multiple accesses to the PCAT1 if these loads reside in different sets of the PCAT1. In the latter case, there needs to be only one access PCAT bit, and one PCAT1 access per cache block. Multiple VPN-PTE entries for multiple problem loads in a cache block can reside in the different ways of a set and thus be obtained in parallel, with the set associativity capping the number of problem loads per cache block for which the PCAT1 can maintain information. The usage of bits of cache block number also simplifies the access of PCAT2 tables and the movement of entries from the PCAT2 to PCAT1, as the same cache block number can be used to access both the L2 cache and the PCAT2, and the entries transferred when the access PCAT bit of the cache block in L2 is set. We briefly evaluate the two options quantitatively in section IV-F.
III-E PCAT/CDTLB/STLB and Page Table Interactions
Like a CDTLB, a PCAT operates like a cache of PTEs, albeit one that is accessed with a load PC rather than a data address. When PTEs are cached, appropriate actions need to be taken to ensure correctness. When an update is made in a PTE, copies cached in CDTLBs must be invalidated using a TLB Shootdown [36, 29]. These changes include the PPN mapping, permissions, and potential updates to the Access bit.
Since the PCAT is accessed with a load PC, and not with a data virtual address, operations using a data virtual address get more involved. Instead of invalidating individual entries using data virtual addresses (as in TLB shootdown) we invalidate the entire PCATs on any change to an entry in the page table (a PCAT Shootdown) as well as the RPA buffer. Since a PCAT entry is created on a (first) CDTLB miss, where the page walker also updates the Access bit in the page table and gives the correct value to the CDTLB, the PCAT will also have the correct value of the Access bit. Any subsequent clearing of the Access bit in the page table (by the OS) results in a PCAT Shootdown. Thus having a PCAT does not impact the operation of the OS page replacement algorithm.
Handling the Dirty bit requires special consideration because the first store to a page may occur after the corresponding entry has already been inserted into a PCAT. Since stores do not consult or update PCAT structures, the Dirty bit recorded in a PCAT entry can become stale even though the Dirty bit in the architectural Page Table and TLB hierarchy has been set. Moreover, a given PTE may appear in the PCATs with different load PCs, further complicating coherence.
A straightforward solution would be to eagerly invalidate the PCATs on any PTE modification. Although correct, this approach would introduce unnecessary invalidations and reduce the performance benefits of PCAX. Instead, we tolerate temporal incoherence in PCAT Dirty bits and rely on the existing TLB hierarchy to ensure correctness.
In a conventional hierarchy, the STLB is checked before invoking a page walk. PCAX extends this mechanism to safely absorb Dirty‑bit incoherence originating from PCATs. Consider the case where the PTE inserted into the CDTLB from PCAT1 has the Dirty bit unset because the PCAT entry was created prior to the first store. That first store has already updated the Dirty bit in the Page Table and propagated it to the STLB and CDTLB (under the standard non-inclusive, non-exclusive rules). If a later load brings an incoherent version of the PTE into the CDTLB from PCAT1, a subsequent store will set the Dirty bit in the CDTLB and attempt to propagate the update toward the STLB. The propagation stops early if the STLB already contains the PTE with the correct Dirty bit set. Thus, correctness is preserved without requiring a page walk. A page walk is only required if the STLB either misses or also contains an unset Dirty bit. At the next periodic PCAT invalidation (Section III-H), the PCAT entries are refreshed with up‑to‑date PTE state, and the above events no longer occur. In our experimental evaluation, with a periodic invalidation interval of 1M instruction, for a small number of benchmarks there are few tens of additional STLB accesses, and a small number (<< 5) additional page walks. Thus, PCAX preserves correctness while minimizing PCAT invalidations through controlled tolerance of temporal Dirty-bit incoherence.
@@ III-F PCAX Coupling with the CDTLB/STLB
An issue is whether PCAT1/PCAT2 (for PCAX) and CDTLB/STLB (for conventional address translation) should be coupled. That is, when a PTE is brought into the PCAT1, as a result of a CDTLB/STLB miss, should the PTE also be placed in the CDTLB/STLB (the structures are coupled), or solely in the PCAT1 (not coupled). Not coupling frees up capacity in the CDTLB/STLB which could be used for other translations whereas not placing a PTE in the CDTLB when it is brought into the PCAT1 results in misses for other proximate load instructions accessing the same PTE. Such instructions cannot get the translation from the PCAT1 (because of different PCs) and it will not be there in the CDTLB.
Our experimental observations suggest that PCAT1/CDTLB coupling is beneficial for most applications. Accordingly, when a PTE is placed in the PCAT1, it is also placed in the CDTLB. For the STLB, given that there is little reason to maintain inclusion between the STLB and the CDTLB/PCATs, we do not couple the PCATs and the STLB. Empirically we have observed that this reduces duplication in the PCAT1/CDTLB and STLB, leading to a slightly lower STLB miss rate.
III-G Replacement in the PCATs
The PCATs can use normal cache replacement policies to determine a victim entry. However, we have found it beneficial to enhance a normal (e.g., LRU) replacement policy with some additional information. To preserve entries that have been deemed to be useful over those that have not been deemed so, we use the Useful bit to exclude an entry from a replacement when possible. Likewise, we have found it beneficial to try to preserve entries that were created as a result of an STLB miss, especially for the PCAT2, since they can potentially provide a translation that might otherwise have to be obtained via an STLB miss.
The STLB Miss (STLBM) bit indicates such entries. The modified LRU replacement algorithm that we employ does not choose an entry with either its Useful or STLBM bit set, unless it doesn’t have a choice. We now describe how the STLBM bit is set. Initially on cold start every entry in the PCATs (and CDTLBs and STLBs) comes from an STLB miss so every entry has its STLBM bit set. However, the PCATs are periodically invalidated, as we discuss in the next section. Invalidation of the PCATs does not involve or impact the CDTLB/STLB, whose contents remain as they are. After an invalidation, when a PCAT entry is created, the STLBM bit is set on not based upon whether the entry came from an STLB miss or hit, with some (not all) entries having their STLBM bit set, and allowing us to use it in the replacement policy.
@@ III-H Periodically Invalidating the PCATs
We periodically invalidate all PCAT structures to prevent entries from persisting indefinitely and causing unproductive behavior. Earlier, we described two scenarios that benefit from periodic invalidation: (1) cold-start behavior affecting the STLBM bit, and (2) temporally incoherent Dirty-bit interactions that can lead to unnecessary page walks. We have also identified a third, subtler pathology that arises from code blocks that become cold.
Consider an instruction cache block A containing code for multiple control-flow paths. Suppose a problem load is detected along one such path, causing a PCAT entry to be created and its Useful bit to be set. If that execution path later becomes cold, the corresponding PCAT2 entry continues to occupy a way in its set. Meanwhile, other problem loads from unrelated code blocks map to the same PCAT2 set and occupy the remaining ways. When block A is reinserted into the L1 instruction cache, its PCAT2 entry is migrated to PCAT1. Importantly, this migration causes the PCAT2 entry to be marked as “accessed” by the local replacement policy—even if the entry is no longer used by the processor while resident in PCAT1. If A is brought into the L1i more frequently than other code blocks that map to the same PCAT2 set, and the PCAT2 replacement policy relies solely on local access information, then the cold entry from A is effectively shielded from eviction. As a result, useful entries belonging to other code blocks are preferentially evicted, while the cold PCAT2 entry persists far longer than desirable. One potential solution is to incorporate actual use information gathered while the entry resides in PCAT1 into the PCAT2 replacement algorithm. However, this approach introduces additional complexity and complicates the replacement logic. Instead, PCAX adopts a simpler and robust strategy: periodically clearing the PCATs. Periodic invalidation flushes stale or cold PCAT entries, allowing the structures to be naturally re-populated with entries corresponding to the currently active working set. This approach effectively eliminates long‑term occupation by cold entries and prevents replacement imbalance. Similar periodic clearing strategies have been shown effective in other microarchitectural structures such as memory dependence predictors—where learned store–load relationships that have gone cold are periodically discarded and relearned [12, 25].
@@ III-I Selectiveness of Loads Using PCAT
PCATs are intended to assist CDTLBs by providing translations that might miss in the CDTLB. PCAT sizes can be kept small because the number of static loads that miss in a CDTLB are small. However, PCAX could also be used for a small number of other ”high value” loads from the set of loads that would normally hit in the CDTLB (whose number is, of course, much larger). Like the data of figure 3 for loads that miss in a CDTLB, we have also found that frequently a dynamic instance of a load that hits in a CDTLB uses the same PTE as the last dynamic instance (data in section IV-D). Judiciously using PCAX for loads that might normally hit in a CDTLB, with reasonably small-sized PCATs, to improve translation latency and energy is left for future work.
@@ IV Evaluation
Processor
@@
L1 DTLB 64-entry, 4-way, 1-cycle, 4-entry MSHR,
2 lookups per cycle, 2 AGU
@@
Page Structure Caches 3-level Split PSC, 2-cycle;
PML4: 2-entry, fully, PDP: 4-entry, fully;
PD: 32-entry, 4-way
@@ IV-A Benchmarks and Experimental Infrastructure
For our evaluation, we use ChampSim [16], a detailed simulator that simulates an out-of-order core with 6-wide fetch and 4-wide issue. Table I summarizes the parameters of the simulated microarchitecture. This version of ChampSim also models a realistic page table walker as found in x86 architectures. Key aspects modeled include: (i) variable latency cost of page table walks and (ii) caching of page walk references in data memory hierarchy as in [40]. Specifically, this models a 4-level page table, a hardware page table walker, and a 3-level split page structured cache. The page walker supports up to 4 concurrent STLB misses. All caches and TLBs use a normal LRU replacement policy. The default page size is 4KB and the default PCAX scheme employs a 64-entry, 2-way PCAT1, a 2K entry, 8-way PCAT2, and a 2-entry RPA Buffer. PCAT1 uses normal LRU while PCAT2 uses the slightly modified LRU replacement which was explained in Section III-G.
For benchmark programs, we use a set of 84 server traces from Qualcomm, which were released after modifications recently [15, 14]. These traces have been widely used in recent papers using the Champsim simulation infrastructure, and studying a variety of issues in instruction supply, memory hierarchies, and microarchitecture [40, 39, 2, 32, 23] and more which can be found in [16].
These traces have a data footprint of 20MB–100MB, which is smaller than those used in other address translation studies [22, 21, 17]. However, the data access patterns produce CDTLB miss rates from 1% to 31%, and CDTLB MPKIs from 1 to 53, with a 64-entry, 4-way associative data TLB. This makes them well suited for evaluating PCAX. Many benchmarks traces have 100M instructions, while some of the traces are smaller (30M instructions) [9]. These traces, released by Qualcomm, have been used in recent works focused on L1 cache and TLB behavior [32, 40, 39, 10].
@@ IV-B Storage Costs
A PCAT entry requires approximately 11.6 bytes: a 20-bit tag, 1 valid bit, 1 useful bit, 1 STLB Miss bit, a 38-bit VPN, and a 4-byte PTE. A 64-entry PCAT1 occupies roughly 744B. PCAT2 configurations with 512, 1024, and 2048 entries require approximately 6KB, 12KB, and 24KB, respectively.
Access PCAT bits are stored per instruction cache block: a 32KB L1 instruction cache with 64-byte blocks requires 512 bits (64B), while a 512KB L2 cache requires 8192 bits (1024B). It would take 2048B with a 1MB L2 cache. The RPA buffer adds an additional 18 bytes.
In our baseline configuration, we use a 64-entry PCAT1 and a 2048-entry PCAT2, resulting in a total storage cost of approximately 25.8 KB. However, as demonstrated in Section IV-G, smaller PCAT2 configurations with 512 and 1024 entries—requiring 7.8 KB and 15.8 KB respectively—achieve similar performance across a range of applications.
@@ IV-C Results Overview
We start our evaluation by presenting some relevant characteristics of the benchmark programs in section IV-D. Then section IV-G evaluates the reduction in MPKIs for primary DTLBs for different PCAT configurations. Section IV-H studies the effect on larger page sizes. Section IV-I presents the service sites for data address translation for a select benchmark set and then sections IV-J and IV-K evaluate the impact on CPI and energy, respectively. Lastly, Section IV-L studies the effect of invalidating the PCATs periodically.
The ordering of the benchmarks on the X-axis in the different figures is different for better visual presentation. Most of our evaluations present results for individual benchmarks. Average results are presented using the geometric mean (GM).
IV-D Benchmark Characteristics
Figure 10 shows the CDTLB Load MPKI for CDTLB sizes of 64, 128, 256, and 1024 entries, with 4-way associativity in each case. For a 64-entry CDTLB, most of the applications have a MPKI >> 7. Even for a (single-level) 1024-entry CDTLB, many applications have a MPKI >> 1.
Figure 10: CDTLB Load MPKI
@@ IV-D Benchmark Characteristics
IV-E PCAT Table Sizes/Hierarchy
@@
Figure 11 shows the total number of problem loads for 64 and 128 entry CDTLBs. The total number is of the order of a few hundred and less than 2000 in almost all cases. These numbers give an idea of how large of a PCAT2 might work well. In some cases, the number of problem loads is higher for the 128-entry CDTLB than for the 64-entry configuration. This occurs in applications with a high MPKI with a 64-entry CDTLB (e.g., 20), which drops significantly (e.g., to 10) with the larger CDTLB. Accounting for 90% of the misses encompasses more problem PCs, which are not included in the 90% when the miss rate is higher.
@@
Since the PCAT1 needs to accommodate only problem loads that are resident in the L1i, we consider this number in Figure 12. The figure presents the maximum number of problem loads resident in the 32KB L1i at any time for 64-entry and 128-entry CDTLBs. This maximum number is in the small hundreds. The average number (not shown) is smaller. As we shall see, PCAT1 sizes that can accommodate a few tens of problem loads (e.g., 64) are adequate in most cases.
Taken together, these observations motivate a hierarchical PCAT organization. Although the cumulative number of problem loads across an application can reach into the thousands—suggesting the need for a reasonably sized PCAT2—only a small subset of these loads is resident in the L1i at any given time. This allows a small, fast PCAT1 to track the active working set of problem loads in the L1i, while a larger PCAT2 retains translations for loads that are evicted from the L1i but likely to be fetched again on an L1i miss.
IV-F PCAT Indexing